Find the value of x for which 2+2x-x2=1
Answer by rothauserc(4718)

multiply both sides of equation by -1 and we have
x2-2x-2=-1, then add 1 to both sides of equation, we get
x2-2x-1=0
x2-2x=1
x(x-1)=1
there is no solution for x